Todrick Company is a merchandiser that reported the following information based on 1,000 units sold: Sales $ 345,000 Beginning merchandise inventory $ 23,000 Purchases $ 230,000 Ending merchandise inventory $ 11,500 Fixed selling expense $ ? Fixed administrative expense $ 13,800 Variable selling expense $ 17,250 Variable administrative expense $ ? Contribution margin $ 69,000 Net operating income $ 20,700 Required: 1. Prepare a contribution format income statement. 2. Prepare a traditional format income statement. 3. Calculate the selling price per unit. 4. Calculate the variable cost per unit. 5. Calculate the contribution margin per unit. 6. Which income statement format (traditional format or contribution format) would be more useful to managers in estimating how net operating income will change in responses to changes in unit sales?
Answer #1
Workings
units 1,000
Sales(A) 345,000
Variable Cost(B=A-C) 276,000
Contribution (C ) 69,000
Fixed Cost (F = C-E) 48,300
Profit (E ) 20,700
Opening Inventory 23,000
Purchases 230,000
Ending Inventory 11,500
Variable production expenses(Opening Inventory +Purchases-Ending Inventory) (G) 241,500
Variable selling and administrative expenses (H=B-G) 34,500
Variable selling expense $ (I) 17,250
Fixed selling expense $ ? (J=F-K) 34,500
Fixed administrative expense $ (K) 13,800
Variable administrative expense $ ? (L=H-I) 17,250
Total Administrative Expenses(K+L) 31,050
Total Distrubution Expenses(I+J) 51,750
1 Traditional Format Income Statement.
Particulars Amount($)
Sales                           345,000.00
Variable production expenses                           241,500.00
Variable selling and administrative expenses                             34,500.00
Contribution                             69,000.00
Fixed selling and administrative expenses                             48,300.00
Profit                             20,700.00
2 Traditional Format Income Statement.
Particulars Amount($)
Revenue                           345,000.00
Cost of sales                           241,500.00
Gross profit                           103,500.00
Administrative expenses                             31,050.00
Distribution expenses                             51,750.00
Research and development expenses                                           -  
Other expenses                                           -  
Total expenses                             82,800.00
Profit before tax                             20,700.00
3 Calculate the selling price per unit Sales Revenue/Units Sold
=345000/1000
345
4 Calculate the variable cost per unit Total Variable Cost / Units Produced
=276000/1000
276
5 Calculate the contribution margin per unit Contribution/ Units sold
=69000/1000
69